You invited 56 people to your graduation party. You can afford to rent 5 tables, round and/or rectangular(each costing the same). Each round table can seat 8 people and each rectangular table can seat 12 people. How many round and rectangular tables should you rent?
....
I just dont know where to start, i'm pretty sure i need to use linear combination.
it would help if you showed me how to do it, it is killing me not knowing waht to do.

thank you very much
hopefully you get it

Answer by scott8148(5895)

let x=round, so 5-x=rectangular

8(x)+12(5-x)=56 ___ 8x+60-12x=56